OpenStudy (anonymous):

Simplify the algebraic expression 4(x + 1) + 7(x + 3) Help please.

OpenStudy (anonymous):

you have to distribute. so you would do 4 times x then 4 times 1 then you would do 7 times x and 7 times 3 them combine like terms

OpenStudy (anonymous):

4*x=4x 4*1=4 7*x=7x 7*3=21 11x+25 is your answer

OpenStudy (anonymous):

4x + 4 + 7x + 21 = 0 Combine like terms. 11x + 25 = 0 To solve for x, move 25 to the other side of the equation. 11x = -25. Divide -25 by 11 to get the value of x.

OpenStudy (anonymous):

@MeganChase Is right !

OpenStudy (anonymous):

Thank you Girls. It 's not been a good couple of days. I might have a few more questions that I would love Y'alls help.

OpenStudy (anonymous):

ok np

OpenStudy (anonymous):

ok...